# Data-Efficient Extremum-Seeking Control Using Kernel-Based Function Approximation

### v1

@@ Abstract
Existing extremum-seeking control (ESC) approaches typically rely on applying repeated perturbations to input parameters and performing measurements of the corresponding performance output. Performing these measurements can be costly in practical applications, e.g., due to the use of resources, making it desirable to reduce the number of performed measurements. Moreover, the required separation between the different timescales in the ESC loop typically results in slow convergence. With these challenges in mind, this work presents an approach aimed at both increasing the convergence rate and reducing the number of measurements that need to be performed. In the proposed approach, input-output data obtained during operation is used to construct online an approximation of the system’s underlying cost function. By using this approximation to perform parameter updates when a decrease in the cost can be guaranteed, instead of performing additional measurements to perform this update, more efficient use is made of the collected data. As a result, reductions in both the required number of measurements and update steps are indeed obtained. In addition, a stability analysis of the novel ESC approach is provided. The benefits of the synergy between kernel-based function approximation and standard ESC is demonstrated in simulation on a multi-input dynamical system.
@@ 1 Introduction
To ensure stable convergence of the input parameters to a small neighborhood of their performance-optimizing values, the different timescales in the ESC loop (the system dynamics, the dither, and optimizer dynamics) should be separated such that the measured performance output remains close to its steady-state solution and the effect of measurement errors on the dynamics of the ESC is small. In classical continuous-time ESC approaches (Krstić and Wang, 2000; Tan et al., 2006, 2010; Scheinker, 2024), for example, this timescale separation is achieved by choosing the frequency of the sinusoidal dither signal sufficiently low compared to the dominant timescale of the system dynamics, and the cut-off frequencies of the high- and low-pass filters used to estimate the gradient of the input-output map via demodulation sufficiently low compared to the dither frequency. However, due to the need for this timescale separation, the convergence rate of the input parameters is typically low. Moreover, the required continuous or repeated application of dither and the required measurements of the corresponding performance output can be costly or undesirable in practical applications, e.g., due to the use of resources such as communication bandwidth or (raw) materials, or due to the transient system responses caused by the dither resulting in scrap or waste products.
@@
To improve the convergence rate, several observer-based ESC approaches have been proposed (Ryan and Speyer, 2010; Gelbert et al., 2012; Guay and Dochain, 2015; Haring and Johansen, 2018). In these approaches, observers are used to obtain local approximations of the input-output map, which provide information about the map (and its derivatives) that is used by the optimization algorithm to adapt the system’s input parameters. The use of observers, instead of the high- and low-pass filters in the classical continuous-time ESC approaches, allows eliminating the timescale associated with the filters and improves the gradient estimate, which enhances the convergence rate of the input parameters. Other proposed approaches instead use a short window of historic data and curve fitting techniques to obtain local approximations of the input-output map (Hunnekens et al., 2014) or the system dynamics (Van Keulen et al., 2020) to estimate the gradient of the input-output map for gradient-based optimization. In Poveda et al. (2021), an approach is proposed that allows the use of both current online measured data and datasets of historic data to obtain instead a non-local approximation of the input-output map whose gradient is used as an estimate of the gradient of the input-output map. Furthermore, besides these observer and approximation-based approaches, in Poveda and Teel (2017), an approach inspired by event-triggered control is proposed in which the convergence rate in a sampled-data ESC setting is increased by sampling the system output as soon as a triggering condition is satisfied, instead of the standard approach in sampled-data ESC of using a fixed waiting time between applying the dither and measuring the corresponding performance output to ensure sufficient timescale separation (Teel and Popović, 2001; Kvaternik and Pavel, 2011; Khong et al., 2013a, b; Hazeleger et al., 2022).
@@
While the aforementioned approaches address the issue of slow convergence in ESC, most of them still require continuous or repeated application of dither and performing measurements of the corresponding system output, which as mentioned before can be costly in practical applications. Exceptions to this are the approaches presented in Hunnekens et al. (2014) and Guay and Dochain (2015). The former approach allows dispensing of the dither signal altogether, while the latter allows dispensing of the dither in some cases. However, the stability proof presented in Hunnekens et al. (2014) is limited to static systems, and in Guay and Dochain (2015) no clear conditions are given for when the dither can be omitted. In Rodrigues et al. (2022, 2023), the issue of the costly nature of performing measurements is partially addressed by using ideas from event-triggered control to determine when to communicate input parameter updates, aimed at reducing the number of updates of the extremum-seeking controller. Using this approach, the use of communication bandwidth is reduced. However, the approach still requires continuous application of dither.
@@
Therefore, in this work, we aim to simultaneously address the issue of slow convergence and reduce the number of input-output measurements that need to be performed. In particular, we adopt a sampled-data ESC setting as in, e.g., Teel and Popović (2001); Hazeleger et al. (2022), since this allows for possible extensions to global optimization algorithms in the future as in Khong et al. (2013a, b) for example. In this setting, we use the input-output data collected during operation of the extremum-seeking controller to construct online an approximation of the steady-state input-output map using kernel-based function approximation. When the approximation is sufficiently accurate in a region around the current optimizer state, it is used to perform a parameter update step without requiring additional inputs to be applied or measurements to be made. In this way, more efficient use is made of previously collected data, which reduces the number of measurements that need to be performed and increases the convergence rate.
@@
In our preliminary work (Weekers et al., 2023), only optimization of static systems maps was considered. The main additional contribution of the current work is therefore the extension of this approach to the more general case of dynamical systems, including a stability proof of the presented approach. Furthermore, we derive novel conditions for assessing whether the kernel-based approximation is sufficiently accurate to be used for a parameter update step. These conditions, which are based on solving convex optimization problems, are less conservative than the closed-form expressions used in Weekers et al. (2023). Finally, we show the benefits of the proposed approach in simulation on a nonlinear dynamical system with multiple inputs.
@@ 2.2 Problem formulation
In the context of ESC, the input $\theta$ and output $y$ of the system $\Sigma_{p}$ can be seen as a vector of tunable system parameters and a to-be-optimized performance variable, respectively. The (constant) input and the output are related in steady-state via the steady-state input-output map $f(\cdot)$ in (3), which thus represents a steady-state cost function for the system $\Sigma_{p}$. However, the system dynamics and consequentially the input-output map $f(\cdot)$ are unknown. The goal in ESC is therefore to solve the steady-state optimization problem
$\theta^{*}:=\underset{\theta\in\mathbb{R}^{n_{\theta}}}{\text{argmin}}\>f(\theta)$ (7)
@@
To solve the steady-state optimization problem (7), typical ESC approaches require the addition of small perturbations $v(\cdot)$ (called dither) to the input parameters $\theta$, and measurements of the corresponding (near steady-state) output $y$. These perturbed inputs and corresponding measured outputs are used by an optimization algorithm aimed at solving (7). Here, we will focus in particular on a class of optimization algorithms that can be described by a difference inclusion (see also, e.g., Teel and Popović (2001) and Hazeleger et al. (2022)). Extensions to a broader class of optimization algorithms, such as in Khong et al. (2013a, b) for example, are left as future work. The class of optimization algorithms that we consider can thus be described by
@@
$Y(\theta):=\begin{bmatrix}f(\theta+v_{1}(\theta))\\ \vdots\\ f(\theta+v_{n_{v}}(\theta))\end{bmatrix}$ (9)
with dither functions $v_{j}(\cdot)$, $j=1,\dots,n_{v}$, and input-output map $f(\cdot)$, maps inputs $\theta$ to vectors containing information regarding (the gradient of) $f(\cdot)$ near $\theta$.
However, as mentioned before, the input-output map $f(\cdot)$ is unknown, and thus the elements of $Y(\theta)$ in (9) cannot be evaluated directly. Instead, they can only be evaluated via measurements of the system output $y$. These output measurements do not exactly match the corresponding values of $f(\cdot)$, due to the system not fully being in steady state at the moment a measurement is taken. Therefore, to analyze convergence properties of the optimization algorithm to the set of minimizers $\mathcal{C}$, and the effect that perturbations to the elements of $Y(\theta)$ in (9) have on these convergence properties, we adopt the following assumptions on the optimization algorithm (8)-(9) and the dither functions $v_{j}(\cdot)$.
@@
    There exists class-$\mathcal{K}_{\infty}$ functions $\omega_{1}$, $\omega_{2}$, and $\rho$, and a nonnegative constant $\delta\in\mathbb{R}_{\geq 0}$, such that, for any $\theta\in\mathbb{R}^{n_{\theta}}$ and $\theta^{*}\in\mathcal{C}$,
    $\omega_{1}(\lVert\theta\rVert_{\mathcal{C}})\leq f(\theta)-f(\theta^{*})\leq\omega_{2}(\lVert\theta\rVert_{\mathcal{C}})$ (10)
@@
    $\underset{\theta^{+}\in F(\theta,Y(\theta))}{\max}\>f(\theta^{+})-f(\theta)\leq-\rho(f(\theta)-f(\theta^{*}))+\delta.$ (11)
@@
    For any $\Delta_{\theta},\Delta_{Y}\in\mathbb{R}_{>0}$, there exists an $L_{Y}\in\mathbb{R}_{>0}$ such that if $\lVert\theta\rVert_{\mathcal{C}}\leq\Delta_{\theta}$ and $\lVert\widetilde{Y}(\theta)\rVert\leq\Delta_{Y}$ with $\widetilde{Y}(\theta)$ a version of $Y(\theta)$ in (9) with inexact (i.e., non-steady-state) measurements of $f(\cdot)$ (cf. (18)), then
@@
    $\lVert v_{j}(\theta)\rVert\leq M_{v}\lVert\theta\rVert_{\mathcal{C}}+c_{v},\quad j=1,\dots,n_{v},$ (13)
@@
Assumption 4 is largely aligned with the assumptions on the optimizer made in Teel and Popović (2001) and Hazeleger et al. (2022) (see also the remark in Khong et al. (2013b, Section 6.2) about the assumptions in Teel and Popović (2001)). The main differences are that: (i) in those works $f(\theta)-f(\theta^{*})$ in (10) and (11) is replaced by a more general locally Lipschitz function $V(\theta)$ (the reason for our stricter assumption will be made clear in Section 4); and (ii) in Hazeleger et al. (2022) additional terms are present in the assumptions that relate to measurable constraints which are not considered here.
Given these assumptions, we recall in Algorithm 1 what we will refer to as a standard parameter update step: a single iteration of the extremum-seeking control algorithm considered in, e.g., Hazeleger et al. (2022) (taking the number of constraints equal to zero) and Teel and Popović (2001). Note that each standard update step described in Algorithm 1 requires $n_{v}$ sequential experiments in which the input-output map $f(\cdot)$ is approximately evaluated for the inputs $\widetilde{\theta}_{i}$ (optimizer state plus dither) via the measured outputs $\widetilde{y}_{i}$. Furthermore, the data collected during these experiments are typically only used for a single algorithm update step, after which they are discarded.
@@
Given the potentially costly nature of performing experiments, we present in the next section an approach for making more efficient use of the collected input-output data $\mathcal{D}_{k}=\{(\widetilde{\theta}_{i},\widetilde{y}_{i}):i=1,\dots,N_{k}\}$. In this approach, the collected data are used to construct online an approximation of the input-output map $f(\cdot)$, which, when it is sufficiently accurate in a sense that will become clear later, is used to determine a search direction and suitable optimizer gain for a gradient-based optimization step without the need for additional experiments.
Algorithm 1 Standard parameter update step
1:Waiting time $T$; number of experiments for a single algorithm update $n_{v}$; dither functions $v_{j}(\cdot)$, $j=1,\dots,n_{v}$; number of experiments performed up to the current algorithm update step $N_{k}$; and optimization algorithm state $\widehat{\theta}_{k}$, time $t_{0,k}$, and system state $x_{0,k}$ at the start of the current algorithm update step $k$.
2:Updated parameters $\widehat{\theta}_{k+1}$.
3:for $i=N_{k}+1,\dots,N_{k}+n_{v}$ do
4:     Apply a constant input including dither $v_{j}(\cdot)$ to the system for a time duration $T$, i.e., $\theta(t)=\widetilde{\theta}_{i}$ $:=\widehat{\theta}_{k}+v_{j}(\widehat{\theta}_{k})$ $\forall t\in[t_{0,k}+(j-1)T,\,t_{0,k}+jT)$ (14) with $j=i-N_{k}$.
5:     Sample the system output $y$ at the end of the time interval, i.e., $\widetilde{y}_{i}:=y(t_{0,k}+jT)=h(\widetilde{x}_{i}(\widehat{\theta}_{k})),$ (15) with $j=i-N_{k}$ and $\widetilde{x}_{i}(\widehat{\theta}_{k}):=x(t_{0,k}+jT,x_{0,k},\widehat{\theta}_{k}+v_{j}(\widehat{\theta}_{k})).$ (16)
6:end for
7:Use the collected samples to update the optimizer state $\widehat{\theta}_{k}$ as in (8)-(9), i.e.,
where $\widehat{\theta}_{k+1}$ can be any element of the set and where
$\widetilde{Y}(\widehat{\theta}_{k}):=\begin{bmatrix}\widetilde{y}_{N_{k}+1}\\ \vdots\\ \widetilde{y}_{N_{k}+n_{v}}\end{bmatrix}=\begin{bmatrix}h(\widetilde{x}_{N_{k}+1}(\widehat{\theta}_{k}))\\ \vdots\\ h(\widetilde{x}_{N_{k}+n_{v}}(\widehat{\theta}_{k}))\end{bmatrix}$ (18)
is an approximation of $Y(\widehat{\theta}_{k})$ with $Y(\cdot)$ as in (9).
8:return $\widehat{\theta}_{k+1}$.
3 Data-efficient extremum-seeking using surrogate modeling

### v2

@@ Abstract
Existing extremum-seeking control (ESC) approaches typically rely on applying repeated perturbations to input parameters and performing measurements of the corresponding performance output. The required separation between the different timescales in the ESC loop means that performing these measurements can be time-consuming. Moreover, performing these measurements can be costly in practice, e.g., due to the use of resources. With these challenges in mind, it is desirable to reduce the number of measurements needed to optimize performance. Therefore, in this work, we present a sampled-data ESC approach aimed at reducing the number of measurements that need to be performed. In the proposed approach, we use input-output data obtained during regular operation of the extremum-seeking controller to construct online an approximation of the system’s underlying cost function. By using this approximation to perform parameter updates when a decrease in the cost can be guaranteed, instead of performing additional measurements to perform this update, we make more efficient use of data collected during regular operation of the extremum-seeking controller. As a result, we indeed obtain a reduction in the required number of measurements to achieve optimization. We provide a stability analysis of the novel sampled-data ESC approach, and demonstrate the benefits of the synergy between kernel-based function approximation and standard ESC in simulation on a multi-input dynamical system.
@@ 1 Introduction
To ensure stable convergence of the input parameters to a small neighborhood of their performance-optimizing values, the different timescales in the ESC loop (the system dynamics, the dither, and optimizer dynamics) should be separated such that the measured performance output remains close to its steady-state solution and the effect of measurement errors on the dynamics of the ESC is small. In classical continuous-time ESC approaches (Krstić and Wang, 2000; Tan et al., 2006, 2010; Scheinker, 2024), for example, this timescale separation is achieved by choosing the frequency of the sinusoidal dither signal sufficiently low compared to the dominant timescale of the system dynamics, and the cut-off frequencies of the high- and low-pass filters used to estimate the gradient of the input-output map via demodulation sufficiently low compared to the dither frequency. However, due to the need for this timescale separation, measuring the (steady-state) performance output can be time-consuming. In addition, it typically causes the convergence rate of the input parameters to be low. Moreover, the required continuous or repeated application of dither and the required measurements of the corresponding performance output can be costly or undesirable in practical applications, e.g., due to the use of resources such as communication bandwidth or (raw) materials, or due to the transient system responses caused by the dither resulting in scrap or waste products.
@@
Many continuous-time approaches have been proposed that improve the convergence rate in ESC. These approaches include Newton-based ESC (Ghaffari et al., 2012) and fixed-time ESC (Poveda and Krstić, 2021) to speed up the learning dynamics, bounded ESC approaches that allow optimization even when the system is unstable (Scheinker and Scheinker, 2016), observer-based approaches (Ryan and Speyer, 2010; Gelbert et al., 2012; Guay and Dochain, 2015; Haring and Johansen, 2018) that improve the gradient estimate and allow eliminating the timescale associated with the high- and low-pass filters in classical approaches, or the use of curve fitting techniques to obtain local (Hunnekens et al., 2014) and non-local (Poveda et al., 2021) approximations of the input-output map for gradient estimation. However, most of these approaches still require timescale separation between the dither and the system dynamics, meaning that obtaining the measurements of the (steady-state) performance output can still be time-consuming. To address this time-consuming nature in a continuous-time ESC context, in Van Keulen et al. (2020) online approximations of the system dynamics are constructed that allow eliminating this separation of timescales between the system dynamics and the dither. Furthermore, in Oliveira et al. (2017), a delay compensation approach is presented for the case where the time-consuming nature stems from a known delay. In a sampled-data context, the time-consuming nature is (partially) addressed in Poveda and Teel (2017) by replacing the typically fixed waiting time between applying inputs and measuring the corresponding steady-state output (see, e.g., Teel and Popović (2001); Kvaternik and Pavel (2011); Khong et al. (2013a, b); Hazeleger et al. (2022)) by a variable waiting time.
@@
In addition, most of the aforementioned approaches still require continuous or repeated application of dither and performing measurements of the corresponding system output, which as mentioned before, can be costly in practice. Exceptions are the approaches presented in Hunnekens et al. (2014) and Guay and Dochain (2015). The former approach allows eliminating the dither signal altogether, while the latter allows eliminating the dither in some cases. However, the stability proof presented in Hunnekens et al. (2014) is limited to static systems, and in Guay and Dochain (2015) no clear conditions are given for when the dither can be omitted. In Rodrigues et al. (2022, 2023), the issue of the costly nature of performing measurements is partially addressed for applications where the costly nature stems from the use of communication bandwidth, by using ideas from event-triggered control. However, this approach does not address the costly nature of performing measurements in a general problem setting.
@@
Therefore, the main contribution of this work is a novel ESC approach which addresses the challenge of both the time-consuming and the costly nature of performing measurements in practical applications, by reducing the number of input-output measurements that need to be performed. In particular, we adopt a sampled-data ESC setting as in, e.g., Teel and Popović (2001); Hazeleger et al. (2022), since this allows for possible extensions to global optimization algorithms as, e.g., in Khong et al. (2013a, b) in the future. In this setting, the novel ESC approach uses the input-output data collected during operation of the extremum-seeking controller to construct online an approximation of the steady-state input-output map using kernel-based function approximation. When the approximation is sufficiently accurate in a region around the current optimizer state, it is used to perform a parameter update step without requiring additional measurements to be made. In this way, more efficient use is made of previously collected data, which reduces the number of measurements that need to be performed. Moreover, unlike the typically many measurements required for offline system identification aimed at model-based optimization, our online approach does not require additional measurement effort to construct the approximation since it uses only the data collected during regular operation of the extremum-seeking controller. A formal stability analysis that shows that our novel ESC approach guarantees (under appropriate tuning conditions) convergence to a neighborhood of the optimum is provided.
@@
In our preliminary work (Weekers et al., 2023), we considered only the optimization of static maps. The additional contribution of the current work is therefore the extension of this approach to the more general case of dynamical systems, including a stability proof of the presented approach. Furthermore, we derive novel conditions for assessing whether the kernel-based approximation is sufficiently accurate to be used for a parameter update step. These conditions, which are based on solving convex optimization problems, are less conservative than the closed-form expressions used in Weekers et al. (2023). Finally, we show the benefits of the proposed approach in simulation on a nonlinear dynamical system with multiple inputs.
@@ 2.2 Problem formulation
In the context of ESC, the input $\theta$ and output $y$ of the system $\Sigma_{p}$ can be seen as a vector of tunable system parameters and a to-be-optimized performance variable, respectively. The (constant) input and the output are related in steady state via the steady-state input-output map $f(\cdot)$ in (3), which thus represents a steady-state cost function for the system $\Sigma_{p}$. However, the system dynamics and consequently the input-output map $f(\cdot)$ are unknown. The goal in ESC is therefore to solve the steady-state optimization problem
$\theta^{*}:=\operatorname*{\arg\min}_{\theta\in\mathbb{R}^{n_{\theta}}}f(\theta)$ (7)
@@
To solve the steady-state optimization problem (7), we will focus in particular on sampled-data ESC approaches in which the optimization algorithm can be described by a difference inclusion (see also, e.g., Teel and Popović (2001) and Hazeleger et al. (2022)). Extensions to a broader class of optimization algorithms, such as in Khong et al. (2013a, b) for example, are left as future work. The class of optimization algorithms that we consider can thus be described by
@@
$Y(\theta):=\begin{bmatrix}f(\theta+v_{1}(\theta))&\cdots&f(\theta+v_{n_{v}}(\theta))\end{bmatrix}^{\intercal}$ (9)
with dither functions $v_{j}(\cdot)$, $j=1,2,\dots,n_{v}$, and input-output map $f(\cdot)$, maps inputs $\theta$ to vectors containing information regarding (the gradient of) $f(\cdot)$ near $\theta$.
However, as mentioned before, the input-output map $f(\cdot)$ is unknown, and thus the elements of $Y(\theta)$ in (9) can only be evaluated via measurements of the system output $y$. We recall how these measurements are obtained during what we will refer to as a standard parameter update step: a single iteration of the ESC algorithm considered in, e.g., Hazeleger et al. (2022) (taking the number of constraints equal to zero) and Teel and Popović (2001). Such an iteration requires $n_{v}$ sequential measurements in which a constant input including dither $v_{j}(\cdot)$ is applied to the system for a time duration $T$, and the output is sampled at the end of the time interval. That is,
$\theta(t)=\widetilde{\theta}_{i}$ $:=\widehat{\theta}_{k}+v_{j}(\widehat{\theta}_{k})$
$\forall t\in[t_{0,k}+(j-1)T,\,t_{0,k}+jT)$ (10)
and
$\widetilde{y}_{i}:=y(t_{0,k}+jT)$ (11)
for each perturbation in the interval with index $j=1,2,\dots,n_{v}$ during the current optimizer iteration $k$. Here, $i=N_{k}+j$ with $N_{k}$ the number of measurements performed up to iteration $k$, and $t_{0,k}$ and $\widehat{\theta}_{k}$ denote, respectively, the time and optimizer state at the start of iteration $k$. Based on the collected samples $\widetilde{y}_{i}$, the optimizer state $\widehat{\theta}_{k}$ is updated as in (8)-(9), i.e.,
where $\widehat{\theta}_{k+1}$ can be any element of the set induced by $F$ and
$\widetilde{Y}(\widehat{\theta}_{k}):=\begin{bmatrix}\widetilde{y}_{N_{k}+1}&\cdots&\widetilde{y}_{N_{k}+n_{v}}\end{bmatrix}^{\intercal}$ (13)
is an approximation of $Y(\widehat{\theta}_{k})$ with $Y(\cdot)$ as in (9), since the system is not fully in steady state at the moment the samples $\widetilde{y}_{N_{k}+1},\widetilde{y}_{N_{k}+2},\dots,\widetilde{y}_{N_{k}+n_{v}}$ are taken. Note that $\widetilde{Y}$ depends on $\widehat{\theta}_{k}$ through the samples of $y$ (cf. (2)).
To analyze convergence properties of the optimization algorithm to the set of minimizers $\mathcal{C}$, and the effect that the discrepancies between $Y(\cdot)$ in (9) and $\widetilde{Y}(\cdot)$ in (13) have on these convergence properties, we adopt the following assumptions on the optimization algorithm (8)-(9) and the dither functions $v_{j}(\cdot)$.
@@
    There exist class-$\mathcal{K}_{\infty}$ functions $\omega_{1}$, $\omega_{2}$, and $\rho$, and a nonnegative constant $\delta\in\mathbb{R}_{\geq 0}$, such that, for any $\theta\in\mathbb{R}^{n_{\theta}}$ and $\theta^{*}\in\mathcal{C}$,
    $\omega_{1}(\lVert\theta\rVert_{\mathcal{C}})\leq V(\theta)\leq\omega_{2}(\lVert\theta\rVert_{\mathcal{C}})$ (14)
@@
    $\underset{\theta^{+}\in F(\theta,Y(\theta))}{\max}\>V(\theta^{+})-V(\theta)\leq-\rho(V(\theta))+\delta$ (15)
    with
    $V(\theta):=f(\theta)-f(\theta^{*}).$ (16)
@@
    For any $\Delta_{\theta},\Delta_{Y}\in\mathbb{R}_{>0}$, there exists an $L_{Y}\in\mathbb{R}_{>0}$ such that if $\lVert\theta\rVert_{\mathcal{C}}\leq\Delta_{\theta}$ and $\lVert\widetilde{Y}(\theta)\rVert\leq\Delta_{Y}$, then
@@
    $\lVert v_{j}(\theta)\rVert\leq M_{v}\lVert\theta\rVert_{\mathcal{C}}+c_{v},\quad j=1,2,\dots,n_{v},$ (18)
@@
Assumption 4 is largely aligned with the assumptions on the optimizer made in Teel and Popović (2001) and Hazeleger et al. (2022) (see also the remark in Khong et al. (2013b, Section 6.2) about the assumptions in Teel and Popović (2001)). The main differences are that: (i) in those works $V(\theta)$ in (14) and (15) is a general locally Lipschitz function $V(\theta)$ instead of being explicitly defined as in (16) (the reason for our stricter assumption will be made clear in Section 4); and (ii) in Hazeleger et al. (2022) additional terms are present in the assumptions that relate to measurable constraints which are not considered here.
@@
Note that each standard update step (10)-(13) requires $n_{v}$ sequential experiments in which the input-output map $f(\cdot)$ is approximately evaluated for the inputs $\widetilde{\theta}_{i}$ (optimizer state plus dither) via the measured outputs $\widetilde{y}_{i}$. Furthermore, the data collected during these experiments are typically only used for a single algorithm update step, after which they are discarded. Given the potentially time-consuming and costly nature of performing these sequential measurements, we present in the next section an approach in which we use the input-output data $\mathcal{D}_{k}=\{(\widetilde{\theta}_{i},\widetilde{y}_{i})\mid i=1,2,\dots,N_{k}\}$ collected during such standard update steps (i.e., during regular operation of the extremum-seeking controller) to construct online an approximation of the input-output map $f(\cdot)$. When this approximation is sufficiently accurate in a sense that we will clarify later, we use it to determine a search direction and suitable optimizer gain for a gradient-based optimization step without the need for additional measurements.
3 Data-efficient extremum-seeking using function approximation